OpenStudy (anonymous):

2x^5 +3x^3 = 2x ? find x?

OpenStudy (shubhamsrg):

whats your progress ?

OpenStudy (anonymous):

what do u mean?

OpenStudy (shubhamsrg):

what have you tried in the question? where are you stuck ?

OpenStudy (anonymous):

kinda confused went it involve the power of 5 , because i tired to let x^3 = u also cant . because there is 2x there

OpenStudy (shubhamsrg):

bring everything to LHS, factor out x, see if it helps ?

OpenStudy (anonymous):

okay then let u = x^2 ? hehe . thanks . kinda new here.

OpenStudy (shubhamsrg):

yep..seems you got it..

OpenStudy (shubhamsrg):

note that since degree of x is 5, you should get 5 roots of the eqn.

OpenStudy (anonymous):

now4 since its divided d ?

OpenStudy (shubhamsrg):

what ?

OpenStudy (anonymous):

the root will be 4 btw thx

OpenStudy (shubhamsrg):

nopes./still there be 5 roots..

OpenStudy (shubhamsrg):

i assume you are missing out that x=0 is also a root..

OpenStudy (anonymous):

how to get x=0??

OpenStudy (shubhamsrg):

by bringing everything on LHS and factoring out x, what do you get ?

OpenStudy (anonymous):

\[\sqrt{2} and \sqrt{\frac{ 1 }{ 2 }}\]

OpenStudy (shubhamsrg):

no no..what i mean is 

OpenStudy (anonymous):

owhhh i see , i cut off those x . haha . thanks :)))

OpenStudy (shubhamsrg):

glad to help.. :)

OpenStudy (anonymous):

r u malaysian?

OpenStudy (shubhamsrg):

nopes..indian..

OpenStudy (anonymous):

owh i c :) thx btw